Can you simplify it?( no negative exponent)
(6x^3y^-4z^5)^-2
Can you find x?
square root of (64^x) = 256
Please help as soon as you can! Thanks

Answer by stanbon(48569)

Can you simplify it?( no negative exponent)
(6x^3y^-4z^5)^-2
--------------
= 1/(36x^6y^-8z^10)
---
= y^8/(36x^6z^10)
========================

Can you find x?
square root of (64^x) = 256
---
(2^6x) = (2^8)
6x = 8
x = 4/3
====================
Cheers,
Stan H.